An aptitude test is designed to measure leadership abilities of the test subjects suppose at the scores on tests are normally distributed with the mean of 550 and a standard deviation of 120 individuals who exceeded 750 on this test are considered to be potential leaders were proportion of the population are considered to be potential leaders run your answered to at least 4 decimal places
Answer #1

Solution :

Given that ,

mean = = 550

standard deviation = = 120

P(x > 750) = 1 - p( x< 750)

=1- p P[(x - ) / < (750 - 550) / 120]

=1- P(z < 1.67)

Using z table,

= 1 - 0.9525

= 0.0475